# Utility-Learning Tension in Self-Modifying Agents

### v1

@@ 7 Algorithmic Self-Modification ($\mathcal{M}_{A}$)
Algorithmic edits cannot cure infinite capacity; on finite capacity, ERM/AERM preserves PAC; and a simple stability meta-policy (cap the step-mass $\sum_{t}\eta_{t}$) controls the generalization gap during algorithmic self-modification.
Figure 2: $\mathcal{M}_{A}$ (Algorithmic axis). Generalization gap (test-train loss) vs. cumulative step–mass $M_{T}=\sum_{t}\eta_{t}$ with a fixed hypothesis class. TwoGate halts updates at a preset budget $B(m)$ and keeps the gap small; Destructive continues updating and exhibits a larger, persistent gap.
No algorithmic cure for infinite VC
[…]
Finite-VC is sufficient with ERM/AERM.
[…]
$R(\hat{h})\ \leq\ \inf_{h\in H}R(h)\ +\ \tilde{O}\!\Big(\sqrt{\tfrac{K+\log(1/\delta)}{m}}\Big).$
Stability meta-policy (step-mass cap).
Assume bounded, Lipschitz, $\beta$-smooth losses (formalized in Appendix 14), and that training examples are sampled uniformly from $S$ during updates. Let a self-modified training run on $H$ use (projected) SGD-like updates with step sizes $\{\eta_{t}\}_{t=1}^{T}$. Define the step-mass $M_{T}:=\sum_{t=1}^{T}\eta_{t}$.
When algorithm changes invalidate ERM assumptions.
If self-modification preserves ERM/AERM over the fixed $H$, Prop. 7 applies directly. If not, we control generalization via algorithmic stability: under standard smooth/Lipschitz conditions, the expected generalization gap scales as $O\!\big((\sum_{t}\eta_{t})/m\big)$; see Thm. 7.
Algorithmic stability via step-mass
Under the conditions above, there exists a constant $C>0$ (problem dependent, independent of $m$) such that
$\mathbb{E}\big[R(\hat{h})-\widehat{R}_{S}(\hat{h})\big]\ \leq\ \frac{C}{m}\,\sum_{t=1}^{T}\eta_{t}\;=\;\frac{C}{m}\,M_{T}.$
In particular, a metacognitive rule that caps $M_{T}\leq B(m)$ guarantees $\mathbb{E}[\mathrm{gap}]=\tilde{O}(B(m)/m)$; choosing $B(m)=\tilde{O}(1)$ yields a $\tilde{O}(1/m)$ gap.
@@ Discussion.
[…]
8 Substrate Self-Modification ($\mathcal{M}_{F}$)
Setting (fixed vs. modifiable).
We analyze substrate edits while holding the specification of $H$, $Z$, and $A$ fixed; switching substrates changes how these are executed (time/space) but not which hypotheses are definable nor which utilities are expressible.
Setting.
The learner runs on a computational substrate $F$ (hardware/VM). A substrate edit is
$F_{t+1}\;=\;\Phi_{F}(F_{t},D_{t},\varphi_{t}).$
Unless otherwise stated, the representation $H$, architecture $Z$, and algorithmic procedure $A$ are held fixed as specifications; switching substrates may change how they are executed (time/space), but not which hypotheses are definable nor which utilities are expressible. PAC learnability is classical (i.i.d., 011 loss unless noted).
@@ 8 Substrate Self-Modification ($\mathcal{M}_{F}$)
(i) Switching among Church–Turing equivalent substrates preserves classical PAC learnability (CT-invariance). (ii) Downgrading to a strictly weaker substrate (e.g., finite-state) can destroy PAC learnability even for problems learnable pre-switch. (iii) “Beyond-CT” substrates do not alter classical PAC guarantees unless they enlarge the induced hypothesis family; if they do, our policy-reachable boundary from §4 applies to the enlarged family.
CT-invariance of PAC learnability
If $F$ and $F^{\prime}$ are Church–Turing equivalent, then a problem that is distribution-freely PAC-learnable when run on $F$ remains distribution-freely PAC-learnable when run on $F^{\prime}$, with the same sample complexity up to constant factors (computation may differ).
Finite-state downgrade can destroy learnability
There exists a binary classification problem that is PAC-learnable on a Turing-equivalent substrate but becomes not distribution-freely PAC-learnable after switching to a fixed finite-state substrate (bounded persistent memory), even when $H$ has finite VC dimension as a specification.
Beyond-CT substrates
If a stronger-than-Turing substrate $F^{\dagger}$ does not enlarge the induced hypothesis family (the measurable set of predictors under consideration), classical PAC learnability is unchanged. If it does enlarge the effective family to $H^{\dagger}$, the learnability boundary is governed by $\sup_{H^{\prime}\in\mathcal{H}^{\dagger}_{\mathrm{reach}}(u)}\text{VC}(H^{\prime})$ exactly as in §4.
@@ Discussion.
Thm. 8 elevates “substrate choice” out of the classical PAC calculus: CT-equivalent machines only affect compute, not sample complexity. Prop. 8 formalizes the “simpleton downgrading” intuition: collapsing persistent memory imposes an information bottleneck that breaks distribution-free guarantees. Prop. 8 shows that any real change to learnability arises only through the induced hypothesis family; our policy-reachable boundary applies verbatim once that family changes. Full proofs appear in Appendix 15.
@@ 9 Outlook
From theory to practice: Why capacity bounds matter.
Practitioners may object that modern deep learning routinely violates PAC sample complexity through implicit regularization—why would capacity bounds matter for self-modifying systems? The answer is sequential compounding risk. A single overparameterized model may generalize well, but a self-modifying agent that repeatedly expands capacity across hundreds of edits accumulates risk that no implicit bias can rescue. Our experiments demonstrate this concretely: the destructive policy reaches test loss 0.409 after unbounded capacity growth, while the capacity-capped policy achieves 0.350—a 17% relative improvement emerging not from a single edit, but from accumulated drift as the reachable family explodes. The Two-Gate policy translates directly into practice: (i) track a capacity proxy $B(\cdot)$ (parameter count for neural architectures, corpus size for retrieval systems, function-library size for tool-using agents); (ii) set a schedule $K(m)\propto\sqrt{m}$ proportional to available training data; (iii) reject edits unless validation improves by margin $\tau$. These checks are computationally cheap and can run in real-time during self-improvement loops. The alternative to capacity bounds is not ”trust implicit regularization”—it is accepting that your system has entered a regime where no learning guarantee is possible.
Multi-axis modification: The realistic frontier.
Real autonomous agents will simultaneously rewrite architectures, swap optimizers, expand tool libraries, and adjust metacognitive policies—not modify a single axis in isolation. Our framework extends naturally: all axes reduce to the same boundary condition on $\sup\mathrm{VC}(\mathcal{H}_{\mathrm{reach}})$. When an agent modifies multiple axes, the induced hypothesis family is determined by the composition of edits—e.g., architecture $Z^{\prime}$ combined with representation $H^{\prime}$ induces $H^{\prime}(Z^{\prime})$, and learnability requires the joint trajectory remain capacity-bounded. This has three practical implications. First, capacity bounds must be enforced globally, not per-axis—a small architectural change combined with representational expansion can yield multiplicative VC growth. Second, axis interactions create emergent capacity explosions that independent per-axis budgets cannot prevent. Third, metacognitive policies become essential: in multi-axis settings, compounding accelerates capacity growth exponentially, making global capacity monitoring the only known mechanism for ex ante safety. The key open challenge is developing compositional capacity proxies that tractably upper-bound $\mathrm{VC}(H^{\prime}(Z^{\prime})\circ\text{tools})$ for complex compositions—the gap between computable bounds $B(\cdot)$ and true VC dimension determines how conservative the Two-Gate policy must be.
@@
The AutoML and NAS communities have achieved remarkable success by treating architecture search as unconstrained optimization. Our results suggest a paradigm shift: future self-modifying systems should ask not ”what maximizes validation accuracy?” but ”what maximizes accuracy subject to capacity remaining PAC-learnable for available data?” This constraint does not eliminate innovation—it channels self-improvement toward sustainable compounding of gains rather than compounding of risk. For open-ended agents operating over months or years, the capacity schedule $K(m)$ can grow with accumulating data $m$, enabling unbounded absolute improvement while maintaining learnability. The failure mode is not self-improvement itself, but uncontrolled self-improvement that outruns the data. Without capacity bounds, seemingly rational modifications can lock in poor performance irreversibly—once an agent crosses into the unbounded regime, no sample complexity guarantees recovery. For high-stakes deployments (medical AI, autonomous vehicles, financial systems), the choice is between principled capacity-aware self-modification and accepting that you have no basis for trust.
@@ 10 Conclusion
We have established a sharp learnability boundary for self-modifying agents: distribution-free PAC guarantees are preserved if and only if the policy-reachable hypothesis family has uniformly bounded capacity. This result unifies representational, architectural, algorithmic, metacognitive, and substrate modifications under a single criterion—the supremum VC dimension of states reachable under the agent’s utility function. The Two-Gate policy provides a computable guardrail that enforces this boundary through validation margins and capacity caps, yielding oracle inequalities at standard VC rates. Our framework reveals that seemingly rational self-modifications can irreversibly destroy learnability when capacity grows without bound, even as they improve immediate performance. As AI systems gain the capability to rewrite their own learning mechanisms, the choice is between principled capacity-aware self-improvement that preserves generalization guarantees, or unconstrained optimization that enters a regime where no learning theory can provide safety assurances. The path to safe open-ended autonomy requires recognizing that self-modification is not a bug to eliminate, but a capability to control—and capacity bounds are the control mechanism that learning theory provides.
@@ 16.3 Bias-Optimal Proof Search
17 Experimental Details
Dataset.
We generate binary labels from a smooth ground-truth score on $X\in\mathbb{R}^{d}$ (Gaussian inputs with mild interactions), passed through a logistic link, plus bounded feature noise and a small independent label-flip rate to induce a nonzero Bayes error. A single draw is split once into train/validation/test and reused across sequential edits.
Representational axis $\mathcal{M}_{H}$.
We simulate representational self-modification by increasing polynomial degree $k=0,1,2,\dots$ in a logistic model. Each candidate is trained on the fixed train set and accepted or rejected by one of four edit policies (Two-Gate, dest_val_nocap, dest_val, dest_train). We report the test 011 loss of the last accepted hypothesis at each step; curves end when an edit is rejected.
Algorithmic axis $\mathcal{M}_{A}$.
We hold the hypothesis class fixed and continue stochastic logistic training while tracking the cumulative step-mass $M_{T}=\sum_{t}\eta_{t}$. We compare a capped TwoGate policy (halt when $M_{T}$ reaches a preset budget) to a destructive policy. We plot the generalization gap (test-train loss) versus $M_{T}$.
17.1 Experiment Hyperparameters
Table 2: Experimental parameters for representational self-modification.
Parameter Value
Train/Val/Test 150/60/1k
Max Degree 30
Noise ($\sigma$) 1.2
Label Flip 35%
L2 Reg. ($C$) 1.0
Capacity ($K$) 31
VC Const. ($c_{0}$) 0.10
Val. Margin ($\tau_{\text{mult}}$) 0.20
Confidence ($\delta_{V}$) 0.05
Seeds 5
Table 3: Experimental parameters for algorithmic self-modification.
Parameter Value
Task Setup Degree 5 logistic
Train Samples 500
Val/Test 1k/2k
Seeds 20
Max Iter. ($T$) 50,000
Learn Rate ($\eta_{0}$) 0.01
Batch Size 32
L2 Reg. $10^{-5}$
Noise ($\sigma$) 0.6
Label Flip 20%
Compared Policies for $\mathcal{M}_{H}$
  1. 1.
    Two-Gate: The safe policy enforcing both a capacity gate and a validation margin.
  2. 2.
    Destructive: Naively accepts any capacity increase if training loss does not increase.
Compared Policies for $\mathcal{M}_{A}$
  1. 1.
    TwoGate: The stability meta-policy. Training is halted when the cumulative step-mass ($M_{T}=\sum\eta_{t}$) exceeds a budget of $B=2.5$.
  2. 2.
    Destructive: The baseline policy. Training proceeds for the full 50,000 iterations with no constraint on step-mass.

### v2

@@ 7 Algorithmic Self-Modification $\mathcal{M}_{A}$
Algorithmic edits cannot cure infinite capacity. On finite capacity, ERM or AERM preserves PAC. When self-modification alters training dynamics beyond ERM assumptions, a simple stability meta-policy based on step-mass controls the generalization gap.
Proposition 9. No algorithmic cure for infinite VC
If $\mathrm{VC}\!\left[H\right]=\infty$, then no distribution-free PAC guarantee is possible for any algorithmic procedure. In particular, algorithmic self-modification cannot restore distribution-free learnability.
Proposition 10. Finite VC is sufficient with ERM or AERM
If $\mathrm{VC}\!\left[H\right]\leq K<\infty$ and the possibly self-modified training procedure is ERM or AERM over $H$, then for any $\delta\in\left]0,1\right[$, with probability at least $1-\delta$ over $S\sim\mathcal{D}^{m}$,
$R\!\left[\hat{h}\right]\ \leq\ \inf_{h\in H}R\!\left[h\right]\ +\ \tilde{O}\!\left[\sqrt{\frac{K+\log\!\left[1/\delta\right]}{m}}\right].$
Stability meta-policy via step-mass.
Assume bounded, Lipschitz, and smooth losses as formalized in Appendix 14, and that training examples are sampled uniformly from $S$ during updates. Let a self-modified training run on $H$ use SGD-like updates with step sizes $\left\{\eta_{t}\right\}_{t=1}^{T}$. Define the step-mass $M_{T}=\sum_{t=1}^{T}\eta_{t}$.
Theorem 11. Algorithmic stability via step-mass
Under the conditions above, there exists a constant $C>0$ independent of $m$ such that
$\mathbb{E}\!\left[R\!\left[\hat{h}\right]-\widehat{R}_{S}\!\left[\hat{h}\right]\right]\ \leq\ \frac{C}{m}\,\sum_{t=1}^{T}\eta_{t}\ =\ \frac{C}{m}\,M_{T}.$
A metacognitive rule that caps $M_{T}\leq B\!\left[m\right]$ guarantees $\mathbb{E}\!\left[\mathrm{gap}\right]=\tilde{O}\!\left[B\!\left[m\right]/m\right]$. Choosing $B\!\left[m\right]=\tilde{O}\!\left[1\right]$ yields a $\tilde{O}\!\left[1/m\right]$ gap.
@@ Discussion.
Proposition 7 says capacity, not optimizer choice, governs distribution-free learnability. Proposition 7 ensures algorithmic edits that continue to output ERM or AERM do not harm PAC guarantees when VC is finite. Theorem 7 offers a simple meta-policy: cap cumulative step-mass to keep the generalization gap small during algorithmic self-modification. Full proofs are in Appendix 14.
8 Substrate Self-Modification $\mathcal{M}_{F}$
Setting, fixed versus modifiable.
We analyze substrate edits while holding the specification of $H$, $Z$, and $A$ fixed. Switching substrates changes how these are executed but not which hypotheses are definable nor which utilities are expressible. PAC learnability is classical in the computable sense unless noted.
A substrate edit is
$F_{t+1}=\Phi_{F}\!\left[F_{t},E_{t},\varphi_{t}\right].$
@@ 8 Substrate Self-Modification $\mathcal{M}_{F}$
First, switching among Church–Turing equivalent substrates preserves classical PAC learnability. Second, downgrading to a strictly weaker substrate, such as finite-state memory, can destroy PAC learnability even for problems learnable pre-switch. Third, stronger-than-Church–Turing substrates do not alter classical PAC guarantees unless they enlarge the induced hypothesis family. If they do, the policy-reachable boundary from Section 4 governs the enlarged family.
Theorem 12. Church–Turing invariance of PAC learnability
If $F$ and $F^{\prime}$ are Church–Turing equivalent, then a problem that is distribution-free PAC learnable when run on $F$ remains distribution-free PAC learnable when run on $F^{\prime}$, with the same sample complexity up to constant factors. Computation may differ.
Proposition 13. Finite-state downgrade can destroy learnability
There exists a binary classification problem that is PAC learnable on a Church–Turing equivalent substrate but becomes not distribution-free PAC learnable after switching to a fixed finite-state substrate with bounded persistent memory, even when $H$ has finite VC dimension as a specification.
Proposition 14. Beyond Church–Turing substrates
If a stronger-than-Church–Turing substrate $F^{\dagger}$ does not enlarge the induced hypothesis family, classical PAC learnability is unchanged. If it enlarges the effective family to $H^{\dagger}$, the learnability boundary is governed by
$\sup_{H^{\prime}\in\mathcal{H}^{\dagger}_{\mathrm{reach}}\!\left[u\right]}\mathrm{VC}\!\left[H^{\prime}\right],$
exactly as in Section 4.
@@ Discussion.
Theorem 8 elevates substrate choice out of the classical PAC calculus: Church–Turing equivalent machines affect compute, not sample complexity. Proposition 8 formalizes that collapsing persistent memory imposes an information bottleneck that breaks distribution-free guarantees. Proposition 8 shows that any real change to learnability arises only through the induced hypothesis family. The policy-reachable boundary applies verbatim once that family changes. Full proofs appear in Appendix 15.
@@ 9 Outlook
From theory to practice, why capacity bounds matter.
Practitioners may object that modern deep learning routinely violates classical worst-case PAC sample complexity through implicit regularization. Why would capacity bounds matter for self-modifying systems? The answer is sequential compounding risk. A single overparameterized model may generalize well, but a self-modifying agent that repeatedly expands capacity across many edits accumulates risk that no implicit bias can reliably rescue.
The Two-Gate policy translates directly into practice: track a capacity proxy $B\!\left[\cdot\right]$, set a schedule $K\!\left[m\right]$ proportional to available data, and reject edits unless validation improves by margin $\tau$. These checks are computationally cheap and can run in real time during self-improvement loops. The alternative to capacity bounds is not trusting implicit regularization. It is accepting that the system has entered a regime where no distribution-free learning guarantee is possible.
Multi-axis modification, the realistic frontier.
Real autonomous agents will simultaneously rewrite architectures, swap optimizers, expand tool libraries, and adjust metacognitive policies. Our framework extends naturally: all axes reduce to the same boundary condition on $\sup\mathrm{VC}\!\left[\mathcal{H}_{\mathrm{reach}}\right]$. When an agent modifies multiple axes, the induced hypothesis family is determined by the composition of edits. Learnability requires the joint trajectory remain capacity-bounded. This has three practical implications. First, capacity bounds must be enforced globally, not per-axis. Second, axis interactions create emergent capacity explosions that independent per-axis budgets cannot prevent. Third, metacognitive policies become essential: in multi-axis settings, compounding accelerates capacity growth, making global capacity monitoring the only known mechanism for ex ante safety.
The key open challenge is developing compositional capacity proxies that tractably upper-bound VC for complex compositions. The gap between computable bounds $B\!\left[\cdot\right]$ and true capacity determines how conservative Two-Gate must be.
@@
AutoML and neural architecture search have achieved success by treating architecture search as unconstrained optimization. Our results suggest a paradigm shift: future self-modifying systems should ask not what maximizes validation accuracy, but what maximizes accuracy subject to capacity remaining PAC learnable for available data. This constraint does not eliminate innovation. It channels self-improvement toward sustainable compounding of gains rather than compounding of risk. For open-ended agents operating over long horizons, the capacity schedule $K\!\left[m\right]$ can grow with accumulating data, enabling unbounded absolute improvement while maintaining learnability. The failure mode is not self-improvement itself, but uncontrolled self-improvement that outruns the data.
@@ 10 Conclusion
We have established a sharp learnability boundary for self-modifying agents: distribution-free PAC guarantees are preserved if and only if the policy-reachable hypothesis family has uniformly bounded capacity. This result unifies representational, architectural, algorithmic, metacognitive, and substrate modifications under a single criterion: the supremum VC dimension of states reachable under the agent’s utility function. The Two-Gate policy provides a computable guardrail that enforces this boundary through validation margins and capacity caps, yielding oracle inequalities at standard VC rates. Our framework reveals that seemingly rational self-modifications can irreversibly destroy learnability when capacity grows without bound, even as they improve immediate performance. As AI systems gain the capability to rewrite their own learning mechanisms, the choice is between principled capacity-aware self-improvement that preserves generalization guarantees, and unconstrained optimization that enters a regime where learning theory provides no safety assurances.